OpenStudy (anonymous):

Susan wants to make pumpkin bread and coffee cakes for the school bake sale. She has 15 eggs and 16 cups of flour in her pantry. Her pumpkin bread recipe uses 2 eggs and 3 cups of flour. Her coffee cake recipe uses 3 eggs and 4 cups of flour. She plans to sell pumpkin bread loaves for $5 each and coffee cake loaves for $4 each. Susan wants to maximize the money raised at the bake sale. Let x represent the number of loaves of pumpkin bread and y represent the number of loaves of coffee cake Susan bakes. What is the objective function for the problem?

OpenStudy (anonymous):

P = 15x + 16y P = 5x + 7y P = 5x + 4y P = 4x + 5y

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

a

OpenStudy (anonymous):

why i need to some one explain why

OpenStudy (anonymous):

hello ??? any one plz

OpenStudy (anonymous):

nvm its b because you have to use 5 cups combined for one and 7 so if you make more than one then thts why the x and y are there i hope u understand what im talking about